A) Suppose A is a 3 x 3 matrix such that det (A) = 1/125 . Find det (5A^-1)


B)Suppose A is a 3 x 3 matrix such that det (A) = 5. Determine det [A^4]
1
Expert's answer
2020-05-19T20:01:43-0400


A)"det(5A^{-1})=5^3det(A^{-1})=\frac{5^3}{det A}=5^6" ;

B)"det(A^4)=(detA)^4=5^4."